OpenStudy (anonymous):

solve for x cos2x=(root3)/2

OpenStudy (saifoo.khan):

x = 15 or 345

OpenStudy (anonymous):

how?

OpenStudy (mathteacher1729):

We can think of it this way: Cos of (some angle) = sqrt(3) / 2 Well, what angles make cos = sqrt(3)/2 ? Now just take those angles, and divide it by 2, since it's really cos (two times some angle...) = sqrt(3)/2 Do you know how to draw a reference triangle?

OpenStudy (anonymous):

so its pi/3 and then divide it by 2?

OpenStudy (anonymous):

\[\cos \theta = \frac{\sqrt{3}}{2}\]for\[\theta = 30^o + 2k(360^o)\]or\[\theta = 330^o+2k(360^o)\]where k is an integer. Take k=0 for principal branch. Then for\[\theta =2x\]\[x=15^o , 165^o\]

OpenStudy (anonymous):

\[\frac{\pi}{12}, \frac{11\pi}{12}\]

myininaya (myininaya):

lokisan where have you been?

OpenStudy (anonymous):

You can obtain these values for the argument of a trig. function from the unit circle.

OpenStudy (saifoo.khan):

lokisan not 165

OpenStudy (mathteacher1729):

http://en.wikipedia.org/wiki/File:Unit_circle_angles_color.svg With color!! Print it, frame it. :D

OpenStudy (anonymous):

\[\cos (2 \times 165^o)=\cos(330^o)=\cos(360^o-30^o)=\cos(30^o)=\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

Hi myininaya, I've been studying/lecturing :)

myininaya (myininaya):

cool!